Suppose eight square have sides of length 1,1,2,3,4,5,6, and 9 inches. Is it possible to piece these tiles together to form a rectangle?
Answer by khwang(438)

The total area:
1+1+2^2+3^2+4^2+5^2+6^2+9^2 = 2+13+ + 41 + 117 = 173
is a prime of oly possiblee dim = 1 x 173 .
And clearly, either side of the final rectangle should be >= 9.
So, it is impossible.
Kenny